OpenStudy (anonymous):

Fan and medal!!!! Indicate the formula for the following conditions. P^c(n, r)

OpenStudy (anonymous):

P(A/B)=P(AnB)/P(B)

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

i really dont understand.. @sherlocked1270

OpenStudy (anonymous):

try googling it

OpenStudy (anonymous):

P(n,r)= n! /(n-r)! C(n,r)= n! / r! (n-r)! P^c(n,r)= P raised to power C(n.r)? (If so, you know C(n,r) ; raise P to power C(n,r)) P(A/B) = P(A ∩ B) / P(B)

OpenStudy (anonymous):

where are yall getting the A and B from?? @ilovefood99 @sherlocked1270

OpenStudy (anonymous):

im not quite sure i looked it up because about 3 weeks ago i had my fiance do these types of problems for me